OpenStudy (anonymous):

Write the equation of a line that passes through the points ( -2, 5) and (-2, 11).

OpenStudy (kkutie7):

first step is to find the slope: \[\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

OpenStudy (kkutie7):

\[\frac{11-5}{-2-2}=-\frac{3}{2}\] now you can write the equation for the line with this: \[y-y_{1}=m(x-x_{1})\] this is were m is the slope.

OpenStudy (kkutie7):

\[y-(5)=(-\frac{3}{2})(x-(-2))\]

OpenStudy (anonymous):

Wrong

OpenStudy (anonymous):

m=11-5/-2-(-2)=6/0

OpenStudy (anonymous):

@not.exactly.a.genius @Kkutie7

OpenStudy (kkutie7):

yes definitely wrong but the set up is right. let me correct that.

OpenStudy (anonymous):

yep

OpenStudy (kkutie7):

\[\frac{11-5}{-2-(-2)}=\frac{6}{0}\] then the line should be straight up and down because the slope is undefined.